# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 7. Additional Complexity Results
We have shown the tractability of resilience for local languages (Theorem 3.3), and shown hardness for four-legged languages (Section 5.2) and for finite reduced languages featuring a word with a repeated letter (Section 6.1). Our results imply a dichotomy over languages with a neutral letter (Section 5.6), but not in general. In this section, we classify some languages not covered by these results, and highlight remaining open cases. We focus on languages that are finite and reduced.
We first illustrate two cases of finite languages that are tractable but not local. One are those we call bipartite chain languages, such as $ab|bc$. For them, we can compute resilience via a variant of the product construction of Theorem 3.3, simply by “reversing” some matches. Another is a class of languages (including for instance $abc|be$) which we show tractable via submodular minimization. These are the only tractable case where we do not know how to reduce to a MinCut problem. We conclude by showing examples of queries not covered by our previous hardness results but for which resilience can be shown intractable, and finally exhibit queries that remain unclassified.
Tractability of Bipartite Chain Languages
@@
Note that chain languages are always finite, because they contain at most $|\Sigma|$ words of length $>2$: indeed, each word of length 3 or more contains intermediate letters which cannot occur in another word. We will focus on chain languages satisfying an extra condition, called being bipartite:
@@
Notice that the definition of BCLs is symmetric: any word $\alpha$ of $L$ can be replaced by the mirror word of $\alpha$ and the resulting language is still a BCL. Also notice that the class of BCLs is incomparable to local languages (even finite local languages): indeed, $ax^{*}b$ and $axb|axc$ are local languages but not BCLs, and Example 7.3 shows BCLs that are not local. Further note that BCLs containing no word of length less than 2 are always reduced. Further, we have the immediate analogue of Section 3.2: the reduced language of a BCL is always a BCL. More generally:
{toappendix}
7.1. Proof of Section 7.1
{lemmarep}
For any BCL $L$, any $L^{\prime}\subseteq L$ is also a BCL.
{toappendix}
We first show that being a chain language is closed under taking subsets:
@@
For any chain language $L$, any $L^{\prime}\subseteq L$ is also a chain language.
Proof.
Let $L^{\prime}\subseteq L$. We check the conditions of chain languages:
  • $L^{\prime}$ cannot contain a word with a repeated letter
  • There cannot be a word $\alpha$ of $L^{\prime}$ sharing an intermediate letter with another word $\beta$ of $L^{\prime}$ without $\alpha$ and $\beta$ being counterexamples for $L$ being a chain language.
This establishes that $L^{\prime}$ is a chain language. ∎
We are now ready to prove Section 7.1:
Proof of Section 7.1.
Let $L$ be a BCL and take $L^{\prime}\subseteq L$. By Lemma 7.4, we know that $L^{\prime}$ is a chain language. Further, if the endpoint graph of $L$ is bipartite, then the same is true of $L^{\prime}$, because the endpoint graph of $L^{\prime}$ is a subgraph of that of $L$. Hence, $L^{\prime}$ is a BCL. ∎
We can show that resilience is tractable for all BCLs, even in combined complexity:
{toappendix}
7.2. Proof of Section 7.2: Bipartite Chain Languages are Tractable
{propositionrep}
[Bipartite chain languages are tractable] If $L$ is a BCL then $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. Moreover, the problem is also tractable in combined complexity: given an $\epsilon$NFA $A$ that recognizes a BCL language $L$, and a database $D$, we can compute $\mathtt{RES}_{\mathrm{bag}}(Q_{L},D)$ in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$.
{proofsketch}
We partition the letters of $\Sigma$ based on the bipartition of the endpoint graph. We then build a flow network like for Theorem 3.3, but we reverse some words depending on the sides of the partition in which their endpoints fall.
{toappendix}
To prove Section 7.2, we first show that an $\epsilon$NFAs recognizing a chain language can be efficiently converted to an explicit representation of the finite language that it recognizes:
@@
Given an $\epsilon$NFA $A$ such that the language $\mathcal{L}(A)$ recognized by $A$ is a chain language, we can compute in time $O(|\Sigma|^{2}\times|A|)$ the finite language $L$ as an explicit list of words.
To prove this, we need an auxiliary claim that intuitively corresponds to the setting where we remove the first and last letters from a chain language:
{claim}
Let $A$ be an $\epsilon$NFA recognizing a finite language where no words contain a repeated letter and where no words share the same letter, i.e., for every word $\alpha\in\mathcal{L}(A)$ all letters of $\alpha$ are distinct, and for any two words $\alpha$ and $\beta$ with $\alpha\neq\beta$ the set of letters of $\alpha$ and of $\beta$ are distinct. Then $\mathcal{L}(A)$ is finite, and we can compute in $O(|A|)$ the finite language $\mathcal{L}(A)$ as an explicit list of words.
Proof.
The finiteness of the language is simply because it has cardinality at most $|\Sigma|+1$: except for the empty word, every word of the language eliminates a letter that cannot be used elsewhere. We assume that $A$ is trimmed (see Definition 3.5); this can be enforced in linear time by Section 3.1.2. We perform two graph traversals to compute in linear time the set $S_{l}$ of states having a path from some initial state via $\epsilon$-transitions, and the set $S_{r}$ of states having a path to some final state via $\epsilon$-transitions.
We first check if $S_{l}\cap S_{r}\neq\emptyset$: if so we add $\epsilon$ to the list of recognized words.
Now, we claim the following (*): for any state $s$ of $A$ which is not in $S_{r}$, there is at most one word $\alpha$ labeling the runs from the initial states of $A$ to $s$. Indeed, assume by way of contradiction that there are two such words $\alpha\neq\beta$. As $A$ is trimmed, let $\gamma$ be a word labeling a run from $s$ to a final state of $A$: as $s\notin S_{r}$, we know that $\gamma$ is non-empty. So $A$ accepts $\alpha\gamma$ and $\beta\gamma$, which are different words because $\alpha\neq\beta$, and which share letters (the ones in $\gamma$), contradicting the assumption on $\mathcal{L}(A)$.
Hence, we compute the list of the non-empty words accepted by $A$ by doing a graph exploration from $S_{l}$, maintaining for each state a word that corresponds to its witness word for the claim (*). The witness word of the states of $S_{l}$ is $\epsilon$. For any other state $s$ that is reached in the traversal, then either it is reached by an $\epsilon$-transition from a state $s^{\prime}$ and its witness word is the same as that of $s^{\prime}$, or it is reached by an $x$-transition for some letter $x$ from a state $s^{\prime}$ and its witness word is $\alpha x$ where $\alpha$ is the witness word of $s^{\prime}$. We store the witness words as pointers to a trie data structure:
  • The root of the trie corresponds to the word $\epsilon$ and the nodes of $S_{l}$ point to the root;
  • If we traverse a transition $s^{\prime}\xrightarrow{\epsilon}s$ then the pointer for $s$ is the same as the pointer for $s^{\prime}$;
@@
When we reach a state of $S_{r}$ then we add a mark the corresponding node in the trie. Note that, as $A$ is trimmed, all leaves of the trie are marked.
At the end of this process, the trie has linear size because the graph traversal is in linear time. Further, les us show that the trie obeys a very strong structural restriction: the least common ancestor of any two marked nodes is the root (**). The claim (**) immediately implies that the trie has the following structure (***): a union of disjoint branches (i.e., the only node with more than one child is the root), with the marked nodes being precisely the set of leaves. Let us now show claim (**). Assume by way of contradiction that there are two marked nodes $n_{1}$ and $n_{2}$ whose lowest common ancestor $n$ is not the root. Then this witnesses that there is a non-empty word $\alpha$ (corresponding to the label of the path from the root to $n$) such that $\alpha\beta$ and $\alpha\gamma$ are accepted by the automaton, for two distinct $\beta$ and $\gamma$ corresponding to paths from $n$ to $n_{1}$ and $n_{2}$ respectively. However, $\alpha\beta$ and $\alpha\gamma$ are two distinct words which share the letters of $\alpha$, a contradiction again.
Thanks to the structure of the trie (***), we can finish by computing explicitly the labels of the paths from the root of the trie to the leaves. Claim (***) implies that this has size linear in the size of the trie, hence the overall process runs in $O(|A|)$. This concludes the proof. ∎
We can now prove our claim on $\epsilon$NFA recognizing chain languages (Lemma 7.5) from Section 7.2:
Proof of Lemma 7.5.
Let $A$ be the $\epsilon$NFA recognizing a language $L\coloneq\mathcal{L}(A)$ which is asserted to be a chain language. We assume that $A$ is trimmed (see Definition 3.5); this can be enforced in linear time by Section 3.1.2.
We first deal in a special way with words of length 00 and $1$ that are accepted by $A$. We first test in linear time as in the proof of Section 7.2 whether $A$ accepts $\epsilon$, and if so, add $\epsilon$ to the list of accepted words.
We then compute in linear time in $A$ the set $\Sigma_{1}$ of letters $a$ such that the single-letter word $a$ is accepted by $L$. We compute these by first computing in linear time the states $S_{l}$ accessible from initial states via $\epsilon$-transitions, computing in linear time the states $S_{r}$ accessible from final states by traversing $\epsilon$-transitions backwards, then going over all transitions with source in $S_{l}$ and target in $S_{r}$ and letting $\Sigma_{1}$ be the set of words which label such transitions. We add all the words of $\Sigma_{1}$ to the list of accepted words.
Hence, after this preprocessing in $O(|A|)$, we can focus on the words of length 2 or more that are accepted by $A$. We now consider every pair of letters $(a,b)\in\Sigma^{2}$ (possibly $a=b$), and compute explicitly the list of words accepted by $A$ that start with $a$ and end with $b$. Let $S_{l,a}$ be the set of states defined in the following way: we take all $a$-labeled transitions whose tail is in $S_{l}$, and let $S_{l,a}$ be the set of all states that are heads of such transitions. Further, let $S_{r,b}$ be the set of states defined in the following way: we take all $b$-labeled transitions whose head in is $S_{r}$, and let $S_{r,b}$ be the set of all states that are tails of such transitions. The sets $S_{l,a}$ and $S_{r,b}$ are computed in time $O(|A|)$. We will now compute explicitly the set of words $\alpha$ such that $a\alpha b\in L$. This clearly amounts to computing the words accepted by the $\epsilon$NFA $A_{a,b}$ obtained from $A$ by changing the initial states to be the states of $S_{l,a}$ and changing the final states to be the states of $S_{r,b}$. Hence, we can compute the list of words in time $O(|A|)$ by Section 7.2. The algorithm has an overall complexity of $O(|A|\times|\Sigma|^{2})$, which concludes the proof. ∎
We can now prove Section 7.2 from Lemma 7.5:
Proof of Section 7.2.
@@
Thanks to Lemma 7.5, we compute in time $O(|A|\times|\Sigma|^{2})$ the language $L\coloneq\mathcal{L}(A)$ as an explicit list of words.
@@
We first process $L$ in linear time to eliminate the words of length 00 or $1$. If $L$ contains $\epsilon$, then resilience is trivial; so in the sequel we assume that $L$ does not contain $\epsilon$. For every letter $a\in\Sigma$ such that $L$ contains the single-letter word $a$, then we know that in the input database $D$ we must remove every $a$-fact for the query to be false; so we can preprocess $D$ to remove all such facts and remove the corresponding one-letter words from $L$. Hence, at the end of this preprocessing, in time $O(|A|+|D|)$, we can assume that $L$ consists of words of length 2 or more.
@@
We now point out that $L$ is reduced. This is by definition of a chain language and from the assumption that the words have length at least 2. Indeed, if $\alpha\in L$ is a strict infix of $\beta\in L$ then $\beta$ must have length at least 3, so there are letters in $\beta$ that are neither the first nor the last letter, by definition of $L$ being a chain language these letters cannot occur in a word of $L\setminus\{\beta\}$, so they do not occur in $\alpha$, so $\alpha$ being an infix of $\beta$ implies that $\alpha$ has length $1$, contradicting the assumption that the words of $L$ have length 2 or more.
@@
  • For every forward word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity +$\infty$+ ∞ from $\mathrm{end}_{a,v,v^{\prime}}$ to $\mathrm{start}_{a,v^{\prime},v^{\prime\prime}}$
  • For every reversed word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity +$\infty$+ ∞ from $\mathrm{end}_{b,v^{\prime},v^{\prime\prime}}$ to $\mathrm{start}_{a,v,v^{\prime}}$.
  • There is an edge with capacity +$\infty$+ ∞ from $t_{\text{source}}$ to every vertex of the form $\mathrm{start}_{a,v,v^{\prime}}$ where $a$ is an endpoint assigned to the source partition, and from every vertex of the form $\mathrm{end}_{a,v,v^{\prime}}$ to $t_{\text{target}}$ where $a$ is an endpoint in the target partition.
@@
We can see that this construction contains a single edge for every fact, whose capacity is the multiplicity of said fact. Every witness of the instance corresponds to a path from source to target by construction. Indeed, for each forward word, we have a path from $t_{source}$ to the edge corresponding to its first letter, then following each letter in succession until we reach the last letter and then $t_{target}$. Further, for each backward word, we have a path from $t_{source}$ to the edge corresponding to its last letter, then following each letter in succession in reverse order, until we reach the first letter and $t_{target}$.

### v3

@@ 7. Additional Complexity Results
We have shown the tractability of resilience for local languages (Theorem 3.13), and shown hardness for four-legged languages (Theorem 5.3) and for finite infix-free languages featuring a word with a repeated letter (Theorem 6.1). Our results imply a dichotomy over languages with a neutral letter (Proposition 5.7), but not in general. In this section, we classify some languages not covered by these results, and highlight remaining open cases. We only consider languages that are infix-free, and mostly study finite languages.
We first illustrate two cases of languages that are tractable but not local. One case are the languages we call bipartite chain languages, such as $ab|bc$, which are finite languages intuitively obtained from a bipartite graph on letter endpoints. For them, we can compute resilience via a variant of the product construction of Theorem 3.13, simply by “reversing” some matches. Another case are the languages we call one-dangling languages, such as $abc|be$, which are intuitively obtained by extending a (possibly infinite) infix-free local language with a single two-letter word with which it shares at most one letter. For them, we can again compute resilience by a more involved encoding to a flow problem. We conclude by showing examples of queries not covered by our previous hardness results but for which resilience can be shown intractable.
7.1. Tractability of Bipartite Chain Languages
@@
Note that chain languages are always finite, because they contain at most $|\Sigma|$ words of length >2>2: indeed, each word of length 3 or more contains intermediate letters which cannot occur in another word. We will focus on chain languages satisfying an extra condition, called being bipartite:
@@
Thus, there are two kinds of chain languages: those that are bipartite, and those which are not. For chain languages $L$ that are not bipartite, we conjecture that $\mathtt{RES}_{\mathrm{set}}(L)$ is always NP-hard. We leave this open, but show hardness for the specific non-bipartite chain language of Example 7.3:
Proposition 7.4.
The resilience problem for the language $ab|bc|ca$ is NP-hard.
Proof.
Figure 13. Completed gadget used in the proof of Proposition 7.4
We claim that the gadget shown in Figure 13 is valid for $ab|bc|ca$. Indeed, the gadget has the following properties: We can see that it meets all the criteria of being a pre-gadget. We can also generate a condensed hypergraph of matches that forms an odd path, and hence this gadget is a valid hardness gadget for $ab|bc|ca$. ∎
@@
We now focus on chain languages that are bipartite, and we will show tractability of resilience for such languages in the rest of this section. Notice that the definition of BCLs is symmetric: any word $\alpha$ of $L$ can be replaced by the mirror word of $\alpha$ and the resulting language is still a BCL. Also notice that the class of BCLs is incomparable to local languages (even finite local languages): indeed, $ax^{*}b$ and $axb|axc$ are local languages but not BCLs, and Example 7.3 shows BCLs that are not local. Further note that BCLs containing no word of length less than 2 are always infix-free. Further, we have the immediate analogue of Lemma 3.14, which we state here and prove in LABEL:apx:redbcl2:
@@
lem:redbcl22 For any BCL $L$, the language $\mathtt{IF}(L)$ is also a BCL.
We can now state our tractability result: the resilience problem is tractable for all BCLs, even in combined complexity and in bag semantics:
Proposition 7.6 (Bipartite chain languages are tractable).
[…]
@@
We can then prove Proposition 7.6 from LABEL:lem:nfachain:
Proof of Proposition 7.6.
@@
Thanks to LABEL:lem:nfachain, we compute in time $O(|A|\times|\Sigma|^{2})$ the language $L\coloneq\mathcal{L}(A)$ as an explicit list of words.
@@
We first process $L$ in linear time to eliminate the words of length 0 or $1$. If $L$ contains $\epsilon$, then resilience is trivial; so in the sequel we assume that $L$ does not contain $\epsilon$. For every letter $a\in\Sigma$ such that $L$ contains the single-letter word $a$, then we know that in the input database $D$ we must remove every $a$-fact for the query to be false; so we can preprocess $D$ to remove all such facts and remove the corresponding one-letter words from $L$. Hence, at the end of this preprocessing, in time $O(|A|+|D|)$, we can assume that $L$ consists of words of length 2 or more.
@@
We now point out that $L$ is infix-free. This is by definition of a chain language and from the assumption that the words have length at least 2. Indeed, if $\alpha\in L$ is a strict infix of $\beta\in L$ then $\beta$ must have length at least 3, so there are letters in $\beta$ that are neither the first nor the last letter, by definition of $L$ being a chain language these letters cannot occur in a word of $L\setminus\{\beta\}$, so they do not occur in $\alpha$, so $\alpha$ being an infix of $\beta$ implies that $\alpha$ has length $1$, contradicting the assumption that the words of $L$ have length 2 or more.
@@
  • For every forward word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity $\infty$ from $\mathrm{end}_{a,v,v^{\prime}}$ to $\mathrm{start}_{b,v^{\prime},v^{\prime\prime}}$
  • For every reversed word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity $\infty$ from $\mathrm{end}_{b,v^{\prime},v^{\prime\prime}}$ to $\mathrm{start}_{a,v,v^{\prime}}$.
  • There is an edge with capacity $\infty$ from $t_{\text{source}}$ to every vertex of the form $\mathrm{start}_{a,v,v^{\prime}}$ where $a$ is an endpoint assigned to the source partition, and from every vertex of the form $\mathrm{end}_{a,v,v^{\prime}}$ to $t_{\text{target}}$ where $a$ is an endpoint in the target partition.
@@
We can see that this construction contains a single edge for every fact, whose capacity is the multiplicity of said fact. Every witness of the instance corresponds to a source-to-target walk by construction. Indeed, for each forward word, we have a path from $t_{source}$ to the edge corresponding to its first letter, then following each letter in succession until we reach the last letter and then $t_{target}$. Further, for each backward word, we have a path from $t_{source}$ to the edge corresponding to its last letter, then following each letter in succession in reverse order, until we reach the first letter and $t_{target}$.